OpenStudy (anonymous):

Input in standard form the equation of the given line. The line that passes through (-1, -3) and (2, 1)

OpenStudy (saifoo.khan):

vijay, u r incorrect for BOOG1's ques.

OpenStudy (saifoo.khan):

1 + 3 ----- 2 + 1 4 --- 3

OpenStudy (vijay):

3y = 4x -7

OpenStudy (saifoo.khan):

y = mx + c 1 = 4(2)/3 + c c = 1 - 8/3 c = -5/3

OpenStudy (saifoo.khan):

3y = 4x - 5

OpenStudy (vijay):

i am correct for boog1's ques

OpenStudy (saifoo.khan):

Nope, those are parellel lines, not over lapping lines.

OpenStudy (vijay):

yeah, they are parallel, but it can take many solutions

OpenStudy (saifoo.khan):

so there wont be any solution.

OpenStudy (anonymous):

saifoo! im giving you medals everytime!

OpenStudy (saifoo.khan):

sorry mlhuck, i typed that in wrong window. lol.

OpenStudy (anonymous):

no worries:) will you help on my next one?

OpenStudy (saifoo.khan):

Sure. :]

OpenStudy (vijay):

y = 3/7 (1-x) and y = 4/7 - 3/7x

OpenStudy (vijay):

when u substitute x values, y can can take many answers

OpenStudy (saifoo.khan):

that dosnt matter, but their slopes are same, & their y intercpts are differnt, tso there wont be any solution.